# XIFBench: Evaluating Large Language Models on Multilingual Instruction Following

### v1

@@ 3.2 Dataset Statistics
Table 1 summarizes XIFBench, which contains 558 instructions and 1,664 evaluation requirements. Each instruction averages 48.1 ± 22.4 tokens, including core instructions and optional input materials. To enable multilingual evaluation, instructions are translated into five languages (Chinese, Russian, Arabic, Hindi, and Swahili), yielding 3,348 instances. The evaluation requirements remain unchanged across languages for consistent evaluation. See Appendix B for details.
@@ 3.3 Evaluation Protocol
Evaluation Framework
We adopt the LLM-based evaluation framework from InfoBench (Qin et al., 2024b), using predefined evaluation requirements to assess how well the responses follow instructions. For each translated instruction and response, we prompt an LLM to generate concise observations and a binary YES/NO decision for each requirement. To ensure cross-linguistic comparability, we provide the original English instruction, evaluation requirements, and metadata (categories and dimensions), keeping them identical across languages as semantic anchors for consistent evaluation. English responses are evaluated on the original English instructions and requirements. See Appendix C for details.
Evaluation Metrics
We define two primary metrics: Requirement Following Rate (RFR) and Instruction Following Rate (IFR). RFR measures the percentage of evaluation requirements correctly satisfied across all instructions, offering a fine-grained view of adherence to specific constraints. IFR quantifies the percentage of instructions where all requirements are met, providing a stricter assessment of overall compliance. These metrics capture LLMs’ ability to follow instructions across constraints and complexities.
Metrics Formulation
Let $\mathcal{I}^{(l)}$ be the set of instructions in language $l\in\{\text{en},\ldots,\text{sw}\}$, each instruction $i^{(l)}\in\mathcal{I}^{(l)}$ has an associated set of evaluation requirements $\mathcal{R}_{i}$, identical across languages. Given an LLM-generated response $o^{(l)}_{i}$ to instruction $i^{(l)}$, the adherence to a requirement $r\in\mathcal{R}_{i}$ is denoted as $e^{(l)}_{i,r}$, a binary value: $1$ if satisfied, and 00 otherwise. The LLM-based evaluation function $\mathcal{E}$ determines requirement adherence:
@@
where $\mathcal{E}$ returns $1$ if the requirement is met, 00 otherwise. English instructions are simplified to:
@@ 4.1 Automatic Evaluation
Table 2: Following rates (%) of different models on XIFBench. The highest rates among closed-source models are in bold, while those among open-source models are underlined. Avg. represents the average rate across all languages.
We employ GPT-4o [^3] We use gpt-4o-2024-08-06 for all evaluations. as the evaluation judge to assess 9 LLMs across two categories: (1) Closed-source LLMs: GPT-4o [^4] We assess gpt-4o-2024-11-20 to avoid self-evaluation., Gemini-2.0-Flash (Google, 2025), and Claude-3.5-Sonnet (Anthropic, 2024); (2) Open-source LLMs: Llama-3.1-70B, Llama-3.1-8B (Meta, 2024), Qwen-2.5-72B, Qwen-2.5-14B, Qwen-2.5-7B (Qwen, 2025), and GLM-4-9B-Chat (GLM, 2024). We access closed-source models via their official APIs, while open-source models run locally using vLLM (Kwon et al., 2023) for efficient inference. We use greedy decoding with a maximum token limit of 4096 to ensure determinism and prevent truncation, keeping all other hyperparameters at default values. LLMs are tested on both sets. We report RFR and IFR scores for each LLM.
@@ 4.2 Agreement Evaluation
Methods En Zh Ru Ar Hi Sw Avg.
DS 70.0 56.7 56.7 55.0 56.7 70.0 60.9
Ours 95.8 96.8 95.1 93.0 95.4 92.3 94.7
Table 3: Agreement rates between GPT-4o and human annotations across different languages (%) within DS and our evaluation protocol. bold means best results.
Figure 3: Cross-lingual RFR performance comparison across constraint categories for three representative models. Each radar chart illustrates the RFR scores across different languages within each constraint category.
Since LLM-based evaluators vary in agreement with humans across languages Hada et al. (2024), we assess the proposed protocol’s reliability through agreement studies. We randomly sample 360 (instruction, response) pairs from six languages, covering both the Easy Set and Hard Set. These pairs are generated by three LLMs with varying scales and sources: Gemini-2.0-Flash, Qwen-2.5-72B, and Glm-4-9B. Other settings follow Section 4.1. We compare our protocol with Direct Scoring (DS) Zheng et al. (2023) whose responses are rated from 1 to 10. Following Jiang et al. (2024); Qin et al. (2024b), responses scoring above 5 are considered to meet all evaluation requirements. Two human annotators assess whether GPT-4o’s evaluation aligns with human judgment for each ground-truth evaluation requirement. GPT-4o also performs Direct scoring for comparison. Finally, we compute agreement rates between GPT-4o and human annotations to assess our protocol’s consistency. See Appendix D.2 for details.
As shown in Table 3, our evaluation protocol outperforms DS across languages, with a slight drop in low-resource languages. The unexpectedly high agreement rate of DS in Swahili does not indicate greater reliability but results from two of the three selected models performing poorly—often scoring below 5 and classified as "not all requirements met"—aligning with human annotations and incidentally boosting agreement.
@@ 5.1 How Do Constraint Categories Affect Cross-lingual Performance?
Figure 4: Cross-lingual IFR performance comparison across instruction complexity levels for three representative models. Each group shows IFR scores across languages per constraint count.
Figure 5: Cross-lingual RFR and IFR performance comparison of culturally universal (CUI) and specific (CSI) instructions across three models. Each group presents RFR and IFR scores per instruction type across languages.
@@ 5.2 How Does Instruction Complexity Impact Cross-lingual Performance?
We use both the Easy Set and Hard Set to obtain instructions ranging from no additional constraints to five constraints. IFR is used to align with instruction-level analysis. Model selection and other experimental settings follow Section 5.1.
@@
Further analysis reveals that higher-resource languages experience lower performance degradation under complex instructions. (1) High-resource languages (English, Chinese) show steady degradation while maintaining acceptable absolute performance. (2) In contrast, mid- and low-resource languages exhibit slightly sharper degradation, resulting in substantial IFR gaps across resource levels as complexity increases.
@@ 6 Conclusion
In this paper, we introduced XIFBench, a benchmark for evaluating the multilingual instruction-following capabilities of LLMs. It features a novel constraint taxonomy, a requirement-based evaluation protocol, and high-quality parallel instruction sets across six languages. Through extensive experiments, we revealed notable variations in LLM performance across resource levels, constraint categories, and instruction complexities. Our findings reveal that LLMs struggle with complex constraints and low-resource languages, highlighting the need for better multilingual training and constraint-aware optimization.
Limitations
While XIFBench provides a comprehensive evaluation framework, certain aspects remain constrained by our design choices: (1) Our benchmark focuses on general-purpose LLMs, excluding reasoning-oriented models (e.g., o1, DeepSeek-R1) due to budget and API constraints. These models, which often infer missing constraints or explore alternative solutions, may exhibit different patterns in constraint adherence, a factor not covered in our current study. (2) As our benchmark relies on parallel translations, cultural accessibility annotations are primarily based on English contexts, which may not fully capture the diversity of cultural factors influencing instruction adherence. Incorporating native speaker annotations could provide a more nuanced perspective. (3) While we identify key factors affecting multilingual instruction following, leveraging these insights to enhance LLMs remains an open challenge. Future research could explore how constraint-based evaluation can inform training or adaptation to improve instruction adherence.
Appendix A Data Construction Details
@@
A.1 Instruction Preparation Details
A.1.1 Seed Instruction Selection
@@
A.1.2 Instruction Filtering and Annotation
@@ 3. If feeling uncertain, you can (recommended from most to least encouraged):
A.2 Constraint Augmentation Details
A.2.1 Constraint Taxonomy
@@
The detailed constraint taxonomy from Section 3.1.2 is presented in Table 6. The examples provided are brainstormed constraints for the instruction Does the 5-second rule exist?, as illustrated in Figure 2.
A.2.2 Constraint Brainstorming
@@ 5. Craft a natural and flowing "combined instruction" by:
A.2.4 Instruction Valiadation
As described in Section 3.1.2, we conducted a manual annotation to evaluate the quality of the combined instructions in the Hard Set. The annotation process followed the guidelines in Table 4, except for the two constraint-related criteria. Two annotators independently assessed all instructions based on three dimensions: clarity, linguistic independence, and cultural accessibility. To measure inter-annotator agreement, we randomly selected 10% of the Hard Set (50 instances) and included them in both annotators’ assignments. Disagreements in this subset were resolved conservatively: if either annotator marked an instruction as defective or culturally specific, it was classified as such. Both annotators were graduate students in computer science and received prior training on the annotation guidelines. We report the proportion of instructions identified as clear, linguistically independent, and culturally universal, along with inter-annotator agreement scores.
@@ A.2.4 Instruction Valiadation
Following the annotation, we removed all unclear and linguistically dependent instructions, along with their corresponding variants in both the Easy Set and the Hard Set, ensuring that each instruction retains all variants across 0 to 5 additional constraints. This process also affected a small number of culturally specific instructions. As a result, the proportion of culturally universal instructions in the Hard Set increased slightly from 70.56% to 70.75%, with a total of 329 culturally universal instructions remaining.
A.3 Requirement Structuring Details
A.3.1 Requirement Decomposition
@@ 5. Is the explanation limited to no more than three sentences?
One might argue that the brainstormed constraints could also serve as evaluation requirements. However, this does not eliminate the necessity of requirement decomposition. The need for decomposition arises not merely from the need to format them as questions, but due to the following reasons: (1) Evaluation requirements are essential for assessing the Easy Set instructions as well as the corresponding components within the Hard Set instructions. (2) The constraints generated by GPT-4o are sometimes compound statements encompassing multiple requirements. For example, a constraint like "Maintain a neutral but informative tone" actually consists of two distinct aspects: neutrality and informativeness. Furthermore, during the constraint combination step, these constraints may be rephrased or semantically altered, making explicit requirement decomposition necessary for accurate evaluation.
A.3.2 Requirement Categorization
@@ A.3.3 Requirement Assessment
Your task is to assess the evaluation requirements of each instruction against specific rules given in the criteria section. We will use evaluation requirements as criteria to assess instruction-following capabilities of Large Language Models (LLMs). These requirements provide fine-grained details on what should be included in model responses. While we will translate the instructions into different languages such as English, Chinese, Russian, Arabic, Hindi, Swahili, and more, the evaluation requirements will remain in English to ensure comparability across languages.
@@
   evaluation_requirements: "1. content requirement - include content: Does the response identify a color that pairs well with teal?"
@@ 3. If feeling uncertain, you can (recommended from most to least encouraged):
Here, we present the details of the requirement assessment introduced in Section 3.1.3. For each evaluation requirement, we assess its quality from four dimensions: explicitness, completeness, atomicity, and categorization. A detailed explanation of each dimension is provided in the annotation guideline (Table 12), resulting in five binary evaluation criteria for each requirement. The categorization dimension is assessed from two perspectives: (1) whether the requirement is correctly categorized and (2) whether it is assigned to the correct dimension. The completeness dimension is evaluated at the instruction level, while the other three dimensions are assessed at the requirement level.
To ensure the reliability of the assessment, two annotators independently annotated a randomly sampled subset, comprising 10% of the instructions from both sets (10 from the Easy Set and 50 from the Hard Set). We calculate both the proportion of high-quality requirements and the inter-annotator agreement on this subset. Disagreements were resolved using a conservative resolution strategy, similar to the approach described in Section A.2.4.
@@
A.4 Multilingual Expansion Details

### v2

@@ 3.2 Dataset Statistics
Table 1 summarizes XIFBench, which contains 558 instructions and 1,664 evaluation requirements. Each instruction averages 48.1 ± 22.4 tokens, including core instructions and optional input materials. To enable multilingual evaluation, instructions are translated into five languages (Chinese, Russian, Arabic, Hindi, and Swahili), yielding 3,348 instances. The evaluation requirements remain unchanged across languages for consistent evaluation. See Appendix D for details.
@@ 3.3 Evaluation Protocol
Evaluation Framework We adopt the widely used LLM-based evaluation protocol from InfoBench [12], using predefined evaluation requirements to assess how well the responses follow instructions. For every instruction and its model-generated response, we prompt an LLM to generate a concise observation and a binary YES/NO decision for each requirement.
While originally designed for English-only or Chinese-only settings, this protocol raises an overlooked question when applied to multilingual evaluation: how does the language of evaluation requirements affect evaluation reliability and consistency? Translating requirements into the same language as the response may seem more natural and contextually aligned, but it may lead to unreliable assessments due to mistranslations and reduce cross-lingual consistency without shared criteria.
To address this, we retain the original English evaluation requirements across all languages, using them as fixed semantic anchors. Specifically, we provide the original English instruction, evaluation requirements, and metadata (categories and dimensions) for each instance, ensuring a shared and stable evaluation framework. This avoids translation-induced inaccuracies and ensures cross-linguistic comparability. See Appendix E for details.
Evaluation Metrics We define two primary metrics: Requirement Following Rate (RFR) and Instruction Following Rate (IFR). RFR measures the percentage of evaluation requirements correctly satisfied across all instructions, offering a fine-grained view of adherence to specific constraints. IFR quantifies the percentage of instructions where all requirements are met, providing a stricter assessment of overall compliance. These metrics capture LLMs’ ability to follow instructions across constraints and complexities.
Metrics Formulation Let $\mathcal{I}^{(l)}$ be the set of instructions in language $l\in\{\text{en},\ldots,\text{sw}\}$, each instruction $i^{(l)}\in\mathcal{I}^{(l)}$ has an associated set of evaluation requirements $\mathcal{R}_{i}$, identical across languages. Given an LLM-generated response $o^{(l)}_{i}$ to instruction $i^{(l)}$, the adherence to a requirement $r\in\mathcal{R}_{i}$ is denoted as $e^{(l)}_{i,r}$, a binary value: $1$ if satisfied, and 0 otherwise. The LLM-based evaluation function $\mathcal{E}$ determines requirement adherence:
@@
where $\mathcal{E}$ returns $1$ if the requirement is met, 0 otherwise. English instructions are simplified to:
@@ 4.1 Automatic Evaluation
Avg. (Open) 89.0 85.8 85.4 77.1 71.8 35.3 74.1 63.1 55.6 55.9 43.1 35.1 10.9 44.0
Table 2: Following rates (%) of different models on XIFBench. The highest rates among closed-source models are in bold, while those among open-source models are underlined. The Avg. column represents the average rate across all languages for each model. The Avg. (Closed) and Avg. (Open) rows represent the average rate for models within their respective categories.
We employ GPT-4o [^2] We use gpt-4o-2024-08-06 for all evaluations. as the evaluation judge to assess 9 LLMs across two categories: (1) Closed-source LLMs: GPT-4o [^3] We assess gpt-4o-2024-11-20 to avoid self-evaluation., Gemini-2.0-Flash [26], and Claude-3.5-Sonnet [27]; (2) Open-source LLMs: Llama-3.1-70B, Llama-3.1-8B [28], Qwen-2.5-72B, Qwen-2.5-14B, Qwen-2.5-7B [29], and GLM-4-9B-Chat [30]. We access closed-source models via their official APIs, while open-source models run locally using vLLM [31] for efficient inference. We use greedy decoding with a maximum token limit of 4096 to ensure determinism and prevent truncation, keeping all other hyperparameters at default values. LLMs are tested on both sets. We report RFR and IFR scores for each LLM.
@@ 4.2 Agreement Evaluation
Considering cross-lingual variation in LLM evaluator agreement [32], we assess our protocol’s reliability and consistency through agreement studies. We compare three settings: (1) Direct Scoring (DS) [10], which holistically rates responses from 1 to 10, considering dimensions such as helpfulness and relevance. Following [11, 12], scores above 5 are considered to meet all evaluation requirements; (2) Ours w/ Trans. Reqs., where evaluation requirements are translated into the instruction’s language, simulating evaluation without a shared semantic anchor to test concerns raised in 3.3; and (3) Ours w/ Eng. Reqs., our proposed protocol using shared English evaluation requirements.
We randomly sample 1080 instruction-response pairs across six languages, covering both the Easy Set and Hard Set. Responses are generated by three LLMs with varying scales: Gemini-2.0-Flash, Qwen-2.5-72B, and Glm-4-9B. GPT-4o performs evaluations for all settings. Two human annotators assess the alignment between GPT-4o’s evaluation and human judgement for each evaluation requirement. We report agreement rates between GPT-4o and human annotations. See Appendix F.2 for details.
Methods En Zh Ru Ar Hi Sw Avg. Std.
DS 71.7 56.7 53.9 55.0 58.3 69.4 60.7 6.5
Ours w/ Trans. Reqs. 93.7 89.1 89.1 86.7 84.6 84.9 88.5 3.1
Ours w/ Eng. Reqs. 95.9 96.7 95.0 93.0 95.5 92.5 94.7 1.6
Table 3: Agreement rates between GPT-4o and human annotations across different languages (%) under DS and our protocol variants. Bold indicates best results.
As shown in Table 3, our protocol (Ours w/ Eng. Reqs.) outperforms DS across languages, with a slight drop in low-resource ones. Its high average agreement (94.7%) and low standard deviation (1.6) confirm its reliability and cross-lingual consistency. The high DS agreement in Swahili doesn’t reflect greater reliability, but rather results from poor model performance–often scoring below 5–coincidentally aligning with human labels. Notably, compared to using translated requirements (Ours w/ Trans. Reqs.), our protocol (Ours w/ Eng. Reqs.) achieves higher agreement across languages (Avg: 94.7% vs. 88.5%), indicating improved reliability. Moreover, it exhibits lower cross-lingual variance (Std: 1.58 vs. 3.10), demonstrating enhanced consistency and fairer comparisons. These results support our initial concern that translating evaluation requirements can compromise both reliability and cross-lingual consistency.
@@ 5.1 How Do Constraint Categories Affect Cross-lingual Performance?
Figure 3: Cross-lingual RFR performance across constraint categories for three representative models. Each radar chart illustrates the RFR scores across different languages within each constraint category.
@@ 5.2 How Does Instruction Complexity Impact Cross-lingual Performance?
Figure 4: Cross-lingual IFR performance across instruction complexity levels for three representative models. Each group shows IFR scores across languages per additional constraint count (+xC).
We use both the Easy Set (no additional constraints) and Hard Set (augmented with 1-5 constraints per instruction) to cover a wide range of instruction complexity. IFR is used to align with instruction-level analysis. Model selection and other experimental settings follow Section 5.1.
@@
Further analysis reveals that performance degradation under complex instructions is not clearly correlated with language resource levels. (1) Across all languages, IFR tends to decline approximately linearly as instruction complexity increases, with relatively similar slopes regardless of resource level. (2) Swahili appears to degrade more steadily in Qwen-2.5-72B and Glm-4-9B, but this is due to its already low baseline performance. Interestingly, we find that degradation patterns vary across models and languages in non-trivial ways. We provide further analysis in Appendix G.3.
@@ 5.3 To What Extent Does Cultural Specificity Influence Instruction Following?
Figure 5: Cross-lingual RFR and IFR performance of culturally universal (CUI) and specific (CSI) instructions across three models. Each group presents following rates per type across languages.
@@ 6 Conclusion
In this paper, we introduced XIFBench, a comprehensive constraint-based benchmark for evaluating the multilingual instruction-following capabilities of LLMs. Featuring a systematic design with constraint-rich instructions across six languages and three methodological innovations for reliable cross-lingual assessment, XIFBench enables fine-grained analysis. Through extensive experiments, we quantified performance disparities across resource levels and provided detailed insights into the influence of language resources, constraint categories, instruction complexity, and cultural specificity on multilingual instruction-following. In summary, XIFBench provides a valuable resource and analytical tool for advancing research and development of robust multilingual LLMs.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (62276077, 62406091, 62350710797, U23B2055), in part by the Guangdong Basic and Applied Basic Research Foundation (2024A1515011205), and in part by the Shenzhen Science and Technology Program (KQTD20240729102154066, ZDSYS20230626091203008).
Appendix A Limitations
While XIFBench provides a comprehensive evaluation framework, we acknowledge several opportunities for future extensions:
  • Our benchmark relies on LLMs for construction and evaluation, supplemented by careful human verification. However, this pipeline may still be susceptible to biases inherent in the "LLM-evaluate-LLM" process. We encourage future work to explore fully manual or alternative human-in-the-loop methodologies for benchmark development.
  • Since our benchmark relies on parallel translations, the cultural accessibility annotations are primarily derived from English contexts, which may not fully capture the cultural diversity across different languages. Directly incorporating annotations from native speakers for non-English instructions would provide a more comprehensive cross-cultural perspective.
  • Our benchmark focuses on general-purpose LLMs, excluding reasoning-specialized models (e.g., o1, DeepSeek-R1, Gemini-2.5-Pro-Preview) due to budget and API constraints. These models, which often infer implicit constraints or explore alternative solutions in their thoughts, may exhibit different patterns in constraint adherence.
  • Another interesting scenario of multilingual instruction-following is code-mixing, i.e., the use of multiple languages in a single instruction. While our benchmark could theoretically evaluate code-mixing by keeping input materials untranslated, we do not include this in our current study. This is because input materials contain diverse content types and would compromise the interpretability of our findings. We leave this as future work.
  • We identify key factors affecting multilingual instruction following, but translating these insights into concrete model improvements remains an open challenge. Future work could explore how our findings can inform training strategies to enhance multilingual instruction following, particularly for semantic-rich constraints and low-resource languages.
Appendix B Broader Impacts
XIFBench is a benchmark designed to evaluate the multilingual instruction-following capabilities of LLMs. By providing a comprehensive evaluation framework, it aims to advance research in this area and promote the development of more robust and capable LLMs. Regarding its broader social impacts, the benchmark can facilitate the development of LLMs that are better equipped to understand and follow instructions in multiple languages, potentially improving accessibility and usability for non-English speakers. This could lead to more inclusive AI systems that cater to a diverse user base. However, a potential negative impact is that developers might prioritize improving model performance on the more widely-spoken languages included in the benchmark, potentially widening the gap between high-resource and low-resource languages.
Appendix C Dataset Construction Details
@@
C.1 Instruction Preparation Details
C.1.1 Seed Instruction Selection
@@
C.1.2 Instruction Filtering and Annotation
@@ 3. If feeling uncertain, you can (recommended from most to least encouraged):
C.2 Constraint Augmentation Details
C.2.1 Constraint Taxonomy
@@
The constraint taxonomy from Section 3.1.2 is presented in Table 6. The examples provided are brainstormed constraints for the instruction Does the 5-second rule exist?, as illustrated in Figure 2.
C.2.2 Constraint Brainstorming
@@ 5. Craft a natural and flowing "combined instruction" by:
C.2.4 Instruction Valiadation
As described in Section 3.1.2, we conducted a manual annotation to evaluate the quality of the combined instructions in the Hard Set. The annotation process followed the guidelines in Table 4, except for the two constraint-related criteria. Two annotators independently assessed all instructions for clarity and linguistic independence, and annotated cultural accessibility. To measure inter-annotator agreement, we randomly selected 10% of the Hard Set (50 instances) and included them in both annotators’ assignments. Disagreements in this subset were resolved conservatively: if either annotator marked an instruction as defective or culturally specific, it was classified as such. Both annotators were graduate students in computer science and received prior training on the annotation guidelines. We report the proportion of instructions identified as clear, linguistically independent, and culturally universal, along with inter-annotator agreement scores.
@@ C.2.4 Instruction Valiadation
Following the annotation, we removed all unclear and linguistically dependent instructions, along with their corresponding variants in both the Easy Set and the Hard Set, ensuring that each instruction retains all variants across 0 to 5 additional constraints. This process also affected a small number of culturally specific instructions. As a result, the proportion of culturally universal instructions in the Hard Set increased slightly from 70.6% to 70.8%, with a total of 329 culturally universal instructions remaining.
Moreover, to assess whether the constraint augmentation process introduces additional culture-specific content, we measured cultural accessibility across both the Easy and Hard Sets. We found that 31.2% of Easy Set instructions and 29.2% of Hard Set instructions are culturally specific. Notably, 97.6% of instructions retain the same cultural accessibility label after augmentation–only 2.4% introduced new culture-specific constraints. This suggests that the constraint combination process has minimal impact on cultural accessibility.
C.3 Requirement Structuring Details
C.3.1 Requirement Decomposition
@@ 5. Is the explanation limited to no more than three sentences?
One might argue that the brainstormed constraints could also serve as evaluation requirements. However, this does not eliminate the necessity of requirement decomposition. The need for decomposition arises not merely from the need to format them as questions, but due to follows: (1) Evaluation requirements are essential for assessing the Easy Set instructions as well as the corresponding components within the Hard Set instructions. (2) The constraints generated by GPT-4o are sometimes compound statements encompassing multiple requirements. For example, a constraint like "Maintain a neutral but informative tone" actually consists of two distinct aspects: neutrality and informativeness. Furthermore, during the constraint combination step, these constraints may be rephrased or semantically altered, making explicit requirement decomposition necessary for accurate evaluation.
C.3.2 Requirement Categorization
@@ C.3.3 Requirement Assessment
Your task is to assess the evaluation requirements of each instruction against specific rules given in the criteria section. We will use these requirements as criteria to assess instruction-following capabilities of LLMs. While we will translate instructions into Chinese, Russian, Arabic, Hindi, and Swahili, evaluation requirements remain in English to ensure comparability.
@@
   evaluation_requirements: "1. content requirement - include content: Does the response identify a color going well with teal?"
@@ 3. If feeling uncertain, you can (recommended from most to least encouraged):
We present the details of the requirement assessment included in Section 3.1.3. For each evaluation requirement, we assess its quality from four dimensions: explicitness, completeness, atomicity, and categorization. A detailed explanation of each dimension is provided in the annotation guideline (Table 12), resulting in five binary evaluation criteria for each requirement. The categorization dimension is assessed from two perspectives: (1) whether the requirement is correctly categorized and (2) whether it is assigned to the correct dimension. The completeness dimension is evaluated at the instruction level, while the other three dimensions are assessed at the requirement level.
To ensure the reliability of the assessment, two annotators independently annotated a randomly sampled subset, comprising 10% of the instructions from both sets (10 from the Easy Set and 50 from the Hard Set). We calculate both the proportion of high-quality requirements and the inter-annotator agreement on this subset. Disagreements were resolved using a conservative resolution strategy, similar to the approach described in Section C.2.4.
@@
C.4 Multilingual Expansion Details